ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 212, 2016

Aaronson, S. and H. Nguyen Near invariance of the hypercube  385

Atzmon, N., D. Ellis and D. Kogan An isoperimetric inequality for conjugation-invariant sets in the symmetric group

139

Barbet, L., M. Dambrine, A. Daniilidis and L. Rifford Sard theorems for Lipschitz functions and applications in optimization

757

Benjamini, I., G. Cohen and I. Shinkar Bi-Lipschitz bijection between the Boolean cube and the Hamming ball

677

Brevik, J. and S. Nollet Grothendieck–Lefschetz theorem with base locus

107

Bu, S. See G. Cai Bucher, M. and A. Talambutsa Exponential growth rates of free

and amalgamated products

521

Cai, G. and S. Bu Periodic solutions of third-order degenerate differential equations in vector-valued functional spaces

163

Chen, E. See D. Zheng Cohen, G. See I. Benjamini Conca, A., E. de Negri and V. Welker A Gorenstein simplicial

complex for symmetric minors

237

Cook, G. C. Bieri–Eckmann criteria for profinite groups  857
Costantini, M. and E. Jabara On locally finite Cpp-groups  123

Dambrine, M. See L. Barbet Daniilidis, A. See L. Barbet de Negri, E. See A. Conca Dettweiler, M. and J. Schmidt Rigid G_{2}-representations and mo-

tives of type G_{2}

81

Ellis, D. See N. Atzmon

Futorny, V. and U. N. Iyer Representations of D_{q}(𝕜[x])  473

Ghasemi, M., S. Kuhlmann and M. Marshall Moment problem in infinitely many variables

989

Greenblatt, M. Applications of an elementary resolution of singularities algorithm to exponential sums and congruences modulo p^{n}

315

Guirardel, V. and G. Levitt Vertex finiteness for splittings of relatively hyperbolic groups 729 Haynes, A. and H. Koivusalo Constructing bounded remainder

sets and cut-and-project sets which are bounded distance to lattices

189

Iyer, U. N. See V. Futorny Jabara, E. See M. Costantini Kaplan, I., N. Levi and S. Shelah The generic pair conjecture for

dependent finite diagrams

259

Karasev, R. N., E. Roldán-Pensado and P. Soberón Measure partitions using hyperplanes with fixed directions

705

Kogan, D. See N. Atzmon Koivusalo, H. See A. Haynes Kuhlmann, S. See M. Ghasemi

Lángi, Z. On a normed version of a Rogers–Shephard type problem  203

Lester, S. and N. Yesha On the distribution of the divisor function and Hecke eigenvalues

443

Levi, N. See I. Kaplan Levitt, G. See V. Guirardel Ma, D. and H. Ren A cycle cover of a 2-edge-connected graph em-

bedded with large face-width on an orientable surface

219

Maillard, P. and E. Paquette Choices and intervals  337

Makarychev, K. and Y. Makarychev Metric extension operators, vertex sparsifiers and Lipschitz extendability

913

Makarychev, Y. See K. Makarychev Marshall, M. See M. Ghasemi Matoušek, J., E. Sedgwick, M. Tancer and U. Wagner Untan-

gling two systems of noncrossing curves

37

Mináč, J., M. Rogelstad, and N. D. Tân Dimensions of Zassenhaus filtration subquotients of some pro-p-groups

825

Moretó, A. and H. N. Nguyen Variations of Landau’s theorem for p-regular and p-singular conjugacy classes

961

Nguyen, H. See S. Aaronson Nguyen, H. N. See A. Moretó

Niemiec, P. Direct integrals of matrices  507

Nollet, S. See J. Brevik Papageorgiou, N. S. and V. D. Rădulescu Nonlinear elliptic

problems with superlinear reaction and parametric concave boundary condition

791

Paquette, E. See P. Maillard Rădulescu, V, D. See N. S. Papageorgiou Ren, H. See D. Ma Rifford, L. See L. Barbet Rogelstad, M. See J. Mináč Roldán-Pensado, E. See R. N. Karasev Sánchez, O. L. On the model companion of partial differential fields

with an automorphism

419

Schmidt, J. See M. Dettweiler Sedgwick, E. See J. Matoušek Shelah, S. See I. Kaplan Shinkar, I. See I. Benjamini Soberón, P. See R. N. Karasev Su, Y. and R. B. Zhang Generalised Jantzen filtration of exceptional

Lie superalgebras

635

Talambutsa, A. See M. Bucher Tân, N. D. See J. Mináč Tancer, M. See J. Matoušek Tikhomirov, K. E. The smallest singular value of random rectangular

matrices with no moment assumptions on entries

289

Wagner, U. See J. Matoušek Welker, V. See A. Conca Xue, H. Fourier–Jacobi periods and the central value of Rankin–

Selberg L-functions

547

Yesha, S. See N. Lester Zhang, R. B. See Y. Su Zheng, D. and E. Chen Bowen entropy for actions of amenable

groups

895

Zweimüller, R. The general asymptotic return-time process  1